point depression data allow us to determine how many moles of particles are dissolved in solution. of 0500 grams of unknown and 20.0 grams of benzene freezes at 2.0O Trand K for benzene (see Table in lab procedure) to find the molality (mol/kg) of the solution, moles, and molar mass of the unknown substance? (Tt for Benzene is 5.5°C and Kf - 5.12C m) A 068 m, 36 6 mol, 0.014 g/mol 068 m 0.014 mol, 36.6 g/mol co 014m, 068 mol, 36.6 g/mol D None of these

plz answer as soon as possible
Answer #1

Δ Tf = 5.5 oC - 2.0 oC = 3.5 oC

use:

Δ Tf = Kf*mb

3.5 = 5.12 *mb

mb= 0.6836 molal

Answer: molality = 0.68 molal

m(solvent)= 20 g

= 2*10^-2 kg

use:

number of mol,

n = Molality * mass of solvent in Kg

= (0.6836 mol/Kg)*(2*10^-2 Kg)

= 1.367*10^-2 mol

Answer: number of mol = 1.4*10^-2 mol

mass(solute)= 0.500 g

use:

number of mol = mass / molar mass

1.367*10^-2 mol = (0.5 g)/molar mass

molar mass = 36.6 g/mol

Answer: molar mass = 36.57 g/mol

Answer:

B